Shift Summary EVE 8/8/99

Collin Wolfe (cwolfe@owlnet.rice.edu)
Mon, 9 Aug 1999 00:00:59 -0500 (CDT)

Shift Summary for EVENING shift, 8 Aug 99

On shift: Collin, Viv
Also present: Jason (Hamm), Theo, Ashkan, Marj, BobT, Valeri, Julie (or her
evil twin), Hogan

Hours of beam available at all intensities: 6.5/8.5
Hours of beam available at nominal intensity: 6.0
Hours of E832 data (special runs not included): 1.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: CsI channel 1137 doesn't like Julie. Also,
problems with crate 51 and the accelerator.

========================================================================

Runs ended this shift:

14150 --- Good E832 data.

========================================================================
Accesses during shift:

Theo and Collin to look at crate 51. Julie, BobT, Jason, Valeri, Hogan
to work on CsI channel 1137 and crate 51.

========================================================================
Overview of shift:

Drift Chambers performed well. Marj chortled with glee over recent
weather patterns. (Humidity dropping means lower DC currents).
CsI channel 1137 dies, crate 51 dies, and the chopper in the accelerator
dies. But other than those problems, it was a smooth hour and a half of
data taking.

========================================================================
Description of Shift:

Hell.

========================================================================

run spill time comment
------------------------------------------------------------------------
14150 15:45 Busy on crate 51 (VV') comes on. Can not be
KFRENDED away. We page Rick. After trying
recycling the downstairs repeater, we realize
we will have to make an access. Theo & Collin
go in.

14150 16:30 Looks bad. Nothing fixes this problem. The
"DATA IN" light is always on in the DYC except
when no modules are attached. On Rick's advice
we decide to swap the DYC. We steal one from
a TRD crate.

17:00 Swapping the DYC has no favorable effect.
We take the crate out of the readout, on
rick's advice. MCR yells at us for taking
too long on access.

17:15 We're out, readout is working (without
ADC crate 8 and crate 51). Val immediately
starts whining about the laser illumination
plot. CsI channel 1137 has dropped out.

17:30 Jason comes in, verifies the channel is dead.

18:00 After conferencing with most of KTeV, we
call in the other half of KTeV to try to
fix channel 1137.

18:15 Worse news. Rosati's pizza oven is broken.

18:30 BobT, Julie, Hogan, Jason go in to swap
the mainframe and divider for channel 1137.
Valeri follows to try to fix crate 51.

18:45 Shift starts looking up as another nearby
Rosati's is located with functional oven
and pizza is ordered.

19:45 Pizza comes (a delightful vegetarian), channel
1137 fixed (swapped divider and mainframe
module), crate 51 fixed (bad discriminator was
feeding signals into crate). All is good.

14151 5 20:20 Data taking starts (spill 5 1st good spill).
SEM around 5E12.

20 20:55 Beam goes away. Problems with the chopper. They
are investigating.

21:40 Beam comes back. SEM around 7E12.

68 21:45 MCR calls to tell us fix is not permanent, and we
will lose beam again in the not too distant future.

91 22:15 Spill aborted

112 22:50 Beam goes away. Must be the downtime the MCR
predicted.

22:56 New message from MCR: Repairing Power Supply
for the Source. No estimate at this time.

23:57 MCR gives estimate for 1/2 hr.

24:00 Hand over to owl shift and get the hell out of
here.

===================================================================

Followup notes on previous shift:

Keep an eye on the e/p for channels 1064,1137,1542.

We did. We shouldn't have.

===================================================================
Items to follow up on next shift:

Keep an eye on the e/p for channels 1064,1137,1542.

If channel 1137 blows its cookies again, just let it go.

Latch monitoring for L1 VV' is not working. Will have to wait
until Tuesday to be fixed.

Do a LASER scan if we lose beam for an hour.